OpenStudy (anonymous):

indefinite integral of 4x^3 sinx^4 dx

zepdrix (zepdrix):

\[\Large \int\limits \sin \color{#F35633}{x^4}\left(4x^3\;dx\right)\]We'll apply a u-substitution. We want to make \(\Large \color{#F35633}{\text{this part}}\) our u. Do you understand how to find du?

OpenStudy (anonymous):

du=(4x^3)dx right? @zepdrix

zepdrix (zepdrix):

Yes, good :)

zepdrix (zepdrix):

Confused on how to apply the substitution? I tried to group the other part in brackets so it'd be easier to line things up.

OpenStudy (anonymous):

so sin(u)du right?? @zepdrix

zepdrix (zepdrix):

\[\Large \int\limits \sin u (du)\]Yes, good! :)

OpenStudy (anonymous):

then I get the derivative of sin(

OpenStudy (anonymous):

then I get the derivative of sin(u) right?? @zepdrix

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

cos(u)du right?? @zepdrix

zepdrix (zepdrix):

When we integrate, the swirly S bar `and` the differential du will both disappear as a part of that process. So when your sin(u) changes, the du should disappear along with the bar. But I think you maybe have your derivatives confused! :O Derivative of sinu = cosu Integral of sinu should give us -cosu, yes? :)

OpenStudy (anonymous):

yes!:) you confused me on the beginning :O @zepdrix

zepdrix (zepdrix):

ah, my bad :3

OpenStudy (anonymous):

ok wait, nvm I get it :p the du no longer exists because it canceled right? :) @zepdrix

zepdrix (zepdrix):

Well, why does this \(\LARGE \int\) no longer exist? :O

OpenStudy (anonymous):

because it goes away with the du right?? @zepdrix

zepdrix (zepdrix):

yes :x I was just a little confused when you said "canceled" lol

OpenStudy (anonymous):

sorry so then I have -cos(u) which its gonna be -cos(x^4) right ?? @zepdrix

zepdrix (zepdrix):

Yay good job \c:/

zepdrix (zepdrix):

We would probably want to throw a +C onto the end to include the entire family of solutions since it's an `indefinite` integral.

OpenStudy (anonymous):

so my answer is -cos(x^4) +c ?? @zepdrix

zepdrix (zepdrix):

Yes! :)

OpenStudy (anonymous):

thank you!! you should help me more often!!!:} @zepdrix